OpenStudy (anonymous):

Set up and simplify an integral that represents the surface obtained by rotating the curve x=sin^2(t), y=sin(3t), 0=

OpenStudy (anonymous):

\[x=\sin ^{2}t\] \[y=\sin(3t)\] \[0 \le t \le (\pi/3)\]